# VersatileFFN: Achieving Parameter Efficiency in LLMs via Adaptive Wide-and-Deep Reuse

### v1

@@ Abstract
[…]
@@ 1 Introduction
The remarkable success of Large Language Models (LLMs) is largely governed by scaling laws, which posit a power-law relationship between model performance and parameter count (kaplan2020scaling; hoffmann2022training). Following this trajectory, the pursuit of state-of-the-art accuracy has led to the proliferation of models with hundreds of billions of parameters. This trend encompasses both dense architectures (touvron2023llama; chen2025pangu; zuo2025man) and sparse Mixture-of-Experts (MoE) variants (jiang2024mixtral; dai2024deepseekmoe; comanici2025gemini; yang2025qwen3; team2025kimi), reflecting a consistent drive toward increasing model scale.
@@
In recent years, a line of research has explored parameter-efficient LLM architectures that maintain high performance under limited memory. Such methods include weight pruning (frantar2023sparsegpt; chen2025pangu), quantization (xiao2023smoothquant; fu2025eaquant), low-rank adaptations (lu2025flrc), and other compression techniques that reduce the storage footprint of pretrained models. However, these approaches primarily aim to approximate the capabilities of the original large model, and thus remain fundamentally constrained by its architectural capacity. They do not augment the model’s representational power, nevertheless, they trade off precision or connectivity for deployability, often hitting an upper bound determined by the base model’s design. There remains a clear need for architectures that are parameter-efficient by design, capable of attaining stronger performance under the same parameter budget, rather than merely compressing existing large models.
In this work, we introduce VersatileFFN, a parameter-efficient architecture that versatilely reuse parameters and computation in feed-forward networks (FFNs), serving as a efficient replacement for the standard FFN in Transformer blocks while leaving the self-attention component unchanged. As illustrated in Figure 1, VersatileFFN integrates two complementary mechanisms: a width-versatile pathway and a depth-versatile pathway, which together enable flexible trade-offs between computational width and depth using a tightly shared parameter set.
@@
Inspired by the dual-process theory of human cognition (kahneman2011thinking), we design an architecture that dynamically allocates computation: “easy” tokens are processed rapidly via a lightweight path, while “hard” tokens receive deeper, iterative reasoning, all under a fixed parameter budget. A difficulty-aware gating mechanism unifies the two pathways, using the expected iteration count from the depth-versatile controller as a proxy for token difficulty. This signal modulates the fusion weights between the outputs of width-versatile FFN and depth-versatile FFN, favoring efficient processing on width-versatile for easy tokens and shifting toward depth-versatile for harder ones. Crucially, both pathways share the same MLP parameters, ensuring added capacity comes from computation, not memory. We conduct extensive experiments across multiple benchmarks and model scales. Models equipped with VersatileFFN consistently outperform other parameters-matched or FLOPs-matched methods, demonstrating the effectiveness of our dual-process design.
@@ 2.1 Mixture-of-Experts
The Mixture-of-Experts (MoE) architecture serves as a foundational framework for scaling model capacity without a proportional increase in computational cost. Originally introduced by (jacobs1991adaptive; jordan1994hierarchical), the concept was later popularized in the large-scale networks by (shazeer2017outrageously). Subsequent advancements, such as GShard (lepikhin2020gshard) and Switch Transformer (fedus2022switch), have demonstrated that replacing standard feed-forward layers in Transformers with MoE layers facilitates efficient pre-training at the trillion-parameter scale, yielding substantial performance improvements. Recent advancements have focused on refining expert granularity and routing strategies (jiang2024mixtral; dai2024deepseekmoe; jin2024moe++; huang2024harder; wang2024remoe), enabling the successful deployment of large-scale MoE models in industrial applications (yang2025qwen3; comanici2025gemini; team2025kimi; zhao2025towards).
Despite their inference efficiency, the massive storage requirements of standard MoE models present significant deployment hurdles. This has catalyzed research into parameter-efficient MoE, utilizing techniques such as expert merging (li2022branch; zhao2024hypermoe), pruning (sarkar2024revisiting; lu2024not), and quantization (nie2022redistribution; dong2024stbllm; huang2024mixture; zhou2025floe). Another emerging direction involves composing lightweight, task-specific modules (e.g. LoRA adapters) into mixture-based systems, as seen in LoRAHub (huang2023lorahub), MoA (feng2024mixture), MoLE (wu2024mixture), and MoRAgent (hanmoragent). While these approaches mitigate some memory constraints, they typically rely on instantiating distinct, physically separate expert parameters, which limits the extent of parameter efficiency compared to architectures designed for intrinsic reuse.
@@ 2.2 Recursive Computation and Adaptive Inference
A promising avenue for maximizing parameter efficiency is the decoupling of model depth from parameter count through recursive computation. Early works such as Universal Transformers (dehghani2018universal) and ALBERT (lan2019albert) demonstrate that cross-layer parameter sharing can induce beneficial inductive biases and improve parameter efficiency for language modeling. Recent theoretical analysis further establishes that such recursive architectures can emulate complex algorithms, acting as universal computers (giannou2023looped; gao2024expressive), and generalize to sequence lengths far beyond those encountered during training (fan2024looped; gong2025makes).
Beyond static parameter sharing, recursive structures facilitate dynamic computation, where the computational budget is adapted to the complexity of the input. Methods like Mixture-of-Depths (raposo2024mixture), Mixture-of-Recurrence (bae2025mixture) and Dynamic resolution network (zhu2021dynamic) dynamically allocate inference FLOPs, allowing models to expend more ”thinking time” on harder tokens or images. This iterative refinement is particularly potent for reasoning tasks, offering significant advantages over static counterparts. Several studies (gatmiry2024can; saunshi2025reasoning; merrill2025little; zhu2025scaling) indicate that increasing computational depth often yields greater performance gains than merely increasing width. While these approaches successfully exploit the depth dimension for efficiency, they typically treat the recurrent layer as a monolithic block, leaving the potential for fine-grained, width-wise parameter reuse largely unexplored, a gap our VersatileFFN aims to bridge.
@@ 3.2 Width-Versatile Mechanism
Construction of Virtual Experts. Rather than allocating discrete weight matrices for each expert, we leverage the shared weights $\mathbf{W}_{proj}$ and $\mathbf{W}_{out}$ as a contiguous parameter substrate. We define a set of $N$ virtual experts by extracting structured, non-overlapping subspaces from the hidden dimension $d_{hidden}$ as illustrated in Figure 2. Specifically, let $d_{expert}$ denote the hidden dimension of a single virtual expert, where $d_{expert}<d_{hidden}$. We employ a strided slicing strategy to map each expert $k\in\{0,\dots,N-1\}$ to a specific view of the shared parameters. This design adheres to two principles: (1) Parameter Efficiency, achieved through the dense reuse of the backbone weights. (2) Functional Orthogonality, ensured by the non-overlapping allocation, which prevents interference between experts.
@@
where $g_{j}$ represents the Softmax-normalized gating probability, and the expert-specific transformation $\mathbf{Y}_{k}(\mathbf{H})$ is:
@@
This sparse activation mechanism decouples computational cost from model capacity. Following previous experience (fedus2022switch), we incorporate an auxiliary load-balancing loss during training to prevent expert collapse.
@@ 3.3 Depth-Versatile Mechanism
Differentiable Loop Prediction. The optimal number of iterations per token is inherently discrete and non-differentiable. To facilitate end-to-end training, we introduce a prediction head $\mathbf{W}_{loop}\in\mathbb{R}^{d\times L_{max}}$ to estimate the required computational depth. The logits for the loop count are computed as $P_{loop}(\mathbf{H})=\mathbf{W}_{loop}\mathbf{H}$. We then employ the Gumbel-Softmax relaxation (jang2016categorical) to sample a differentiable probability vector $\mathbf{p}\in\Delta^{L_{max}-1}$:
@@
where $\mathbf{g}\sim\text{Gumbel}(0,1)$ and $\tau$ is the temperature parameter.
@@ 3.4 Difficulty-Aware Fusion
15:   Compute final output $\mathbf{Y}$ by fusing pathways via Eq. 3
@@ 4.1 Implementation Details
Experiments Setup. To validate our approach, we conducted pre-training experiments on the FineWeb-Edu dataset (penedo2024fineweb), our implementation is based on OLMo2 (olmo20242). Specifically, we train a 354M parameter model on 40B tokens and a 720M parameter model on 70B tokens, where both token sets are subsets derived from a larger 100B token corpus (huggingfacefw_2024). Following the architectural settings of the open-source OLMo2 models, we set the hidden dimensions to 1,024 for 354M model and 1,536 for the 720M model. Both models consist of 15 layers and utilize SwiGLU for activation and RMSNorm (zhang2019root) for normalization. Besides, The OLMo2 tokenizer with a vocabulary size of 50,280 is employed in our models.
@@
We primaly conduct experiments under a continued pre-training setting. Specifically, we initialize the model using the pre-trained checkpoint and continue training for one epoch on the original corpus. For all models, we set the training sequence length to 4,096 tokens. AdamW optimizer is employed with a weight decay of 0.1 and a gradient clipping threshold of 1.0. The learning rate follows a cosine decay schedule, beginning with a warm-up phase for the first 5% of total steps and subsequently decaying to 10% of the peak value. Regarding model-specific hyperparameters, we use a global batch size of 1.5M tokens with a peak learning rate of $5\times 10^{-4}$ for the 354M model, and a global batch size of 1M tokens with a peak learning rate of $4\times 10^{-4}$ for the 720M model. Furthermore, for ours and traditional MoE, we incorporate an auxiliary load balancing loss with a weighting coefficient of $1\times 10^{-5}$.
Evaluation Benchmark Setup. We evaluate our method on a comprehensive set of benchmarks: PIQA (Bisk2020), HellaSwag (zellers-etal-2019-hellaswag), OpenBookQA (OBQA) (OpenBookQA2018), SciQ (Welbl2017CrowdsourcingMC), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (clark2018think), CommonsenseQA (COMM)  (clark-etal-2019-boolq), and Winogrande (WINO) (sakaguchi2019winogrande). We utilize OLMES (olmo20242) to assess the performance under zero-shot setting.
@@
Base Model Parameters: 354.71M
@@
Base Model Parameters: 720.81M
@@
Method Avg. Params FFN FLOPs
(M) (M)
Base Model Parameters: 354.71M
Base 47.98 354.71 377.49
MoE 51.48 543.59 471.86
2-Loop 51.47 354.71 754.98
4-Loop 51.98 354.71 1509.96
6-Loop 51.94 354.71 2264.96
VersatileFFN 52.33 354.90 1236.08
Base Model Parameters: 720.81M
Base 53.83 720.81 849.35
MoE 55.87 1145.69 1061.69
2-Loop 55.83 720.81 1698.70
4-Loop 56.33 720.81 3397.40
6-Loop 56.55 720.81 5096.10
VersatileFFN 57.03 721.09 2586.38
Versatile Setup of Width and Depth. To determine the optimal configuration for the VersatileFFN architecture, specifically the number of active experts and total experts in the width (i.e., $k$ and $N$ in Eq. 7) and the maximum loops in the depth (i.e., $L_{\max}$ in Eq. 11), we conduct ablation studies using the 300M model trained on 40B tokens. In these experiments, we maintained one shared expert for the width-versatile pathway, while disabling virtual experts within the depth-versatile pathway. The training loss and average accuracy across the above eight academic benchmarks during training is illustrated in Figure 3. For width-versatile, the losses (and average accuracies) for varying expert configurations, specifically 4-choose-1, 8-choose-2, 8-choose-4, and 16-choose-4 are 2.655 (50.16%), 2.633 (50.98%), 2.633 (50.96%), and 2.634 (50.22%), respectively. We select the 8-choose-2 setting as our primary experimental configuration due to its higher performance. For depth-versatile, when the number of loops is 1, 2, 4, and 6, the final training losses are 2.654, 2.631, 2.625, and 2.623, respectively. The average accuracies are 50.09%, 51.47%, 51.98%, and 51.94%. It can be seen that 4 loops can achieve the best balance between accuracy and computation. Therefore, we set the maximum number of loops to 4 in our main experiment.

### v2

@@ Abstract
The rapid scaling of Large Language Models (LLMs) has achieved remarkable performance, but it also leads to prohibitive memory costs. Existing parameter‑efficient approaches such as pruning and quantization mainly compress pretrained models without enhancing architectural capacity, thereby hitting the representational ceiling of the base model. In this work, we propose VersatileFFN, a novel feed‑forward network (FFN) that enables flexible reuse of parameters in both width and depth dimensions within a fixed parameter budget. Inspired by the dual‑process theory of cognition, VersatileFFN comprises two adaptive pathways: a width‑versatile path that generates a mixture of sub‑experts from a single shared FFN, mimicking sparse expert routing without increasing parameters, and a depth‑versatile path that recursively applies the same FFN to emulate deeper processing for complex tokens. A difficulty‑aware gating dynamically balances the two pathways, steering “easy” tokens through the efficient width‑wise route and allocating deeper iterative refinement to “hard” tokens. Crucially, both pathways reuse the same parameters, so all additional capacity comes from computation rather than memory. Experiments across diverse benchmarks and model scales demonstrate the effectiveness of the method. The code is available at https://github.com/huawei-noah/noah-research/tree/master/VersatileFFN.
@@ 1 Introduction
The remarkable success of Large Language Models (LLMs) is largely governed by scaling laws, which posit a power-law relationship between model performance and parameter count (Kaplan et al., 2020; Hoffmann et al., 2022). Following this trajectory, the pursuit of state-of-the-art accuracy has led to the proliferation of models with hundreds of billions of parameters. This trend encompasses both dense architectures (Touvron et al., 2023; Chen et al., 2025; Zuo et al., 2025) and sparse Mixture-of-Experts (MoE) variants (Jiang et al., 2024; Dai et al., 2024; Comanici et al., 2025; Yang et al., 2025; Team et al., 2025), reflecting a consistent drive toward increasing model scale.
@@
In recent years, a line of research has explored parameter-efficient LLM architectures that maintain high performance under limited memory. Such methods include weight pruning (Frantar and Alistarh, 2023; Chen et al., 2025), quantization (Xiao et al., 2023; Fu et al., 2025), low-rank adaptations (Lu et al., 2025), and other compression techniques that reduce the storage footprint of pretrained models. However, these approaches primarily aim to approximate the capabilities of the original large model, and thus remain fundamentally constrained by its architectural capacity. They do not augment the model’s representational power, nevertheless, they trade off precision or connectivity for deployability, often hitting an upper bound determined by the base model’s design. There remains a clear need for architectures that are parameter-efficient by design, capable of attaining stronger performance under the same parameter budget, rather than merely compressing existing large models.
In this work, we introduce VersatileFFN, a parameter-efficient architecture that versatilely reuse parameters and computation in feed-forward networks (FFNs), serving as an efficient replacement for the standard FFN in Transformer blocks while leaving the self-attention component unchanged. As illustrated in Figure 1, VersatileFFN integrates two complementary mechanisms: a width-versatile pathway and a depth-versatile pathway, which together enable flexible trade-offs between computational width and depth using a tightly shared parameter set.
@@
Inspired by the dual-process theory of human cognition (Kahneman, 2011), we design an architecture that dynamically allocates computation: “easy” tokens are processed rapidly via a lightweight path, while “hard” tokens receive deeper, iterative reasoning, all under a fixed parameter budget. A difficulty-aware gating mechanism unifies the two pathways, using the expected iteration count from the depth-versatile controller as a proxy for token difficulty. This signal modulates the fusion weights between the outputs of width-versatile FFN and depth-versatile FFN, favoring efficient processing on width-versatile for easy tokens and shifting toward depth-versatile for harder ones. Crucially, both pathways share the same MLP parameters, ensuring added capacity comes from computation, not memory. We conduct extensive experiments across multiple benchmarks and model scales. Models equipped with VersatileFFN consistently outperform other parameters-matched or FLOPs-matched methods, demonstrating the effectiveness of our dual-process design.
@@ 2.1 Mixture-of-Experts
The Mixture-of-Experts (MoE) architecture serves as a foundational framework for scaling model capacity without a proportional increase in computational cost. Originally introduced by (Jacobs et al., 1991; Jordan and Jacobs, 1994), the concept was later popularized in the large-scale networks by (Shazeer et al., 2017). Subsequent advancements, such as GShard (Lepikhin et al., 2020) and Switch Transformer (Fedus et al., 2022), have demonstrated that replacing standard feed-forward layers in Transformers with MoE layers facilitates efficient pre-training at the trillion-parameter scale, yielding substantial performance improvements. Recent advancements have focused on refining expert granularity and routing strategies (Jiang et al., 2024; Dai et al., 2024; Jin et al., 2024; Huang et al., 2024a; Wang et al., 2024), enabling the successful deployment of large-scale MoE models in industrial applications (Yang et al., 2025; Comanici et al., 2025; Team et al., 2025; Zhao et al., 2025).
Despite their inference efficiency, the massive storage requirements of standard MoE models present significant deployment hurdles. This has catalyzed research into parameter-efficient MoE, utilizing techniques such as expert merging (Li et al., 2022; Zhao et al., 2024), pruning (Sarkar et al., 2024; Lu et al., 2024), and quantization (Nie et al., 2022; Dong et al., 2024; Huang et al., 2024b; Zhou et al., 2025). Another emerging direction involves composing lightweight, task-specific modules (e.g. LoRA adapters) into mixture-based systems, as seen in LoRAHub (Huang et al., 2023), MoA (Feng et al., 2024), MoLE (Wu et al., 2024), and MoRAgent (Han et al., 2025). While these approaches mitigate some memory constraints, they typically rely on instantiating distinct, physically separate expert parameters, which limits the extent of parameter efficiency compared to architectures designed for intrinsic reuse.
@@ 2.2 Recursive Computation and Adaptive Inference
A promising avenue for maximizing parameter efficiency is the decoupling of model depth from parameter count through recursive computation. Early works such as Universal Transformers (Dehghani et al., 2018) and ALBERT (Lan et al., 2019) demonstrate that cross-layer parameter sharing can induce beneficial inductive biases and improve parameter efficiency for language modeling. Recent theoretical analysis further establishes that such recursive architectures can emulate complex algorithms, acting as universal computers (Giannou et al., 2023; Gao et al., 2024), and generalize to sequence lengths far beyond those encountered during training (Fan et al., 2024; Gong et al., 2025).
Beyond the paradigm of static parameter sharing, recursive structures facilitate dynamic computation, where the computational budget is adapted to the complexity of the input. Recent advancements such as Mixture-of-Depths (Raposo et al., 2024), Mixture-of-Recurrence (Bae et al., 2025) and Dynamic resolution network (Zhu et al., 2021) dynamically allocate inference FLOPs, allowing models to expend more ”thinking time” on harder tokens or images. This iterative refinement is particularly potent for reasoning tasks, offering significant advantages over static counterparts. Several studies (Gatmiry et al., 2024; Saunshi et al., 2025; Merrill and Sabharwal, 2025; Zhu et al., 2025) indicate that increasing computational depth often yields greater performance gains than merely increasing width. In addition, HRM (Wang et al., 2025) and TRM (Jolicoeur-Martineau, 2025) show that iteratively reasoning models can outperform larger static counterparts. While these approaches successfully exploit the depth dimension for efficiency, they typically treat the recurrent layer as a monolithic block, leaving the potential for fine-grained, width-wise parameter reuse largely unexplored. Ours aims to bridge this gap by introducing a multi-dimensional sharing mechanism that exploits redundancy across both depth and width.
@@ 3.2 Width-Versatile Mechanism
Construction of Virtual Experts. Rather than allocating discrete weight matrices for each expert, we leverage the shared weights $\mathbf{W}_{proj}$ and $\mathbf{W}_{out}$ as a contiguous parameter substrate. We define a set of $N$ virtual experts by extracting structured, non-overlapping subspaces from the hidden dimension $d_{hidden}$ as illustrated in Figure 2. Specifically, let $d_{expert}$ denote the hidden dimension of a single virtual expert. To strictly satisfy the non-overlapping requirement, we impose the constraint $N\times d_{expert}\leq d_{hidden}$, ensuring that the total capacity of the virtual experts does not exceed the shared parameter space. We employ a strided slicing strategy to map each expert $k\in\{0,\dots,N-1\}$ to a specific view of the shared parameters. This design adheres to two principles: (1) Parameter Efficiency, achieved through the dense reuse of the backbone weights. (2) Functional Orthogonality, ensured by the non-overlapping allocation, which prevents interference between experts.
@@
where $g_{k}$ represents the Softmax-normalized gating probability, and the expert-specific transformation $\mathbf{Y}_{k}(\mathbf{H})$ is:
@@
This sparse activation mechanism decouples computational cost from model capacity. Following previous experience (Fedus et al., 2022), we incorporate an auxiliary load-balancing loss during training to prevent expert collapse.
@@ 3.3 Depth-Versatile Mechanism
Differentiable Loop Prediction. The optimal number of iterations per token is inherently discrete and non-differentiable. To facilitate end-to-end training, we introduce a prediction head $\mathbf{W}_{loop}\in\mathbb{R}^{d\times L_{max}}$ to estimate the required computational depth. The logits for the loop count are computed as $P_{loop}(\mathbf{H})=\mathbf{W}_{loop}\mathbf{H}$. We then employ the Gumbel-Softmax relaxation (Jang et al., 2016) to sample a differentiable probability vector $\mathbf{p}\in\Delta^{L_{max}-1}$:
@@
where $\mathbf{g}\sim\text{Gumbel}(0,1)$ and $\tau$ is the temperature.
@@ 3.4 Difficulty-Aware Fusion
15:   Compute output $\mathbf{Y}$ by fusing pathways via Eq. 3
@@ 4.1 Implementation Details
Experiments Setup. To validate our approach, we conducted pre-training experiments on the FineWeb-Edu dataset (Penedo et al., 2024), our implementation is based on OLMo2 (OLMo et al., 2024). Specifically, we train a 354M parameter model on 40B tokens, a 720M parameter model on 70B tokens and a 1.21B parameter model on 100B tokens. The 40B and 70B token sets are subsets derived from the 100B token corpus (HuggingFaceFW, 2024). Following the architectural settings of the open-source OLMo2 models, we set the hidden dimensions to 1,024 for 354M model, 1,536 for the 720M model and 2,048 for 1.21B model. All models consist of 15 layers and utilize SwiGLU for activation and RMSNorm (Zhang and Sennrich, 2019) for normalization. Besides, The OLMo2 tokenizer with a vocabulary size of 50,280 is employed in our models. Further training setting and evaluation benchmarks are provided in Appendix B.
@@
Base Model Params: 354.71M
@@
Base Model Params: 720.81M
@@
Base Model Params: 1.21B
Base 2.385 73.07 54.18 38.00 88.90 69.12 35.12 40.05 55.49 56.74
MoE 2.278 74.21 59.37 41.00 91.10 70.35 41.14 41.20 58.80 59.65
2-Loop 2.328 74.59 59.08 39.60 91.10 71.93 41.47 40.70 57.38 59.48
4-Loop 2.322 73.97 59.41 40.70 91.50 72.86 41.51 42.86 57.87 60.09
6-Loop 2.320 74.76 59.86 41.20 91.40 71.58 40.47 41.93 59.19 60.05
VersatileFFN 2.319 74.65 60.17 40.80 91.90 73.16 41.14 43.73 58.17 60.47
@@
Method Avg. Params (B) FLOPs (G)
Base 47.98 0.35 0.38
MoE 51.48 0.54 0.47
2-Loop 51.47 0.35 0.75
4-Loop 51.98 0.35 1.50
6-Loop 51.94 0.35 2.25
VersatileFFN 52.33 0.35 1.24
Base 53.83 0.72 0.85
MoE 55.87 1.15 1.06
2-Loop 55.83 0.72 1.70
4-Loop 56.33 0.72 3.40
6-Loop 56.55 0.72 5.10
VersatileFFN 57.03 0.72 2.59
Base 56.74 1.21 1.51
MoE 59.65 1.97 1.89
2-Loop 59.48 1.21 3.02
4-Loop 60.09 1.21 6.04
6-Loop 60.05 1.21 9.06
VersatileFFN 60.47 1.21 4.69
Versatile Setup of Width and Depth. To determine the optimal configuration for the VersatileFFN architecture, specifically the number of active experts and total experts in the width (i.e., $k$ and $N$ in Eq. 7) and the maximum loops in the depth (i.e., $L_{\max}$ in Eq. 11), we conduct ablation studies using the 354M model trained on 40B tokens. In these experiments, we maintained one shared expert for the width-versatile pathway, while disabling virtual experts within the depth-versatile pathway. The training loss and average accuracy across eight academic benchmarks during training is illustrated in Figure 3. For width-versatile, the losses (and average accuracies) for varying expert configurations, specifically 4-choose-1, 8-choose-2, 8-choose-4, and 16-choose-4 are 2.655 (50.16%), 2.633 (50.98%), 2.633 (50.96%), and 2.634 (50.22%), respectively. We select the 8-choose-2 setting as our primary experimental configuration due to its higher performance. For depth-versatile, when the number of loops is 1, 2, 4, and 6, the final training losses are 2.654, 2.631, 2.625, and 2.623, respectively. The average accuracies are 50.09%, 51.47%, 51.98%, and 51.94%. It can be seen that 4 loops can achieve the best balance between accuracy and computation. Therefore, we set the maximum number of loops to 4 in our main experiment.